OpenStudy (a1234):

The equation r = (sqrt A)/(pi) gives the radius r of circle with area A. What is the radius of a circle with the given area? Write your answer as a simplified radical and as a decimal rounded to the nearest hundredth. a. 50 ft^2

OpenStudy (jdoe0001):

\(\bf r=\sqrt{\cfrac{{\color{red}{ Area}}}{\pi}}\qquad {\color{red}{ Area}}=50\ ft^2\implies r=\sqrt{\cfrac{{\color{red}{ 50}}}{\pi}} \)

OpenStudy (a1234):

I think we can simplify that further.